[Brahms-l] Shift report 20050503 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue May 03 2005 - 15:54:09 EDT
Shifter leader was FR
Shifters were 



 - There was no beam during this shift. Machine development and then
   equipment failures. Current status "Machine setup".
                                                                                
 - Only pedestal runs: 14405, 14406
                                                                                
 - TPC's and Drift chambers are in park mode.
   All magnets set in StandBy.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 3 15:54:50 2005

This archive was generated by hypermail 2.1.8 : Tue May 03 2005 - 15:55:19 EDT